# Polymatroid Schubert varieties

### v1

@@ 4. Polymatroid Schubert varieties
Proof of 1.4.
Let $S=\mathbb{Q}[y_{\mathbf{s}}:\text{$\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$}]$, graded by $\deg y_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. Using the notation of 4.14, define a map
@@
where $\mathbf{b}$ is any basis of $\mathbf{s}$. The image of $y_{\mathbf{s}}$ does not depend on $\mathbf{b}$ by 4.14.
@@
$y_{\mathbf{s}}y_{\mathbf{s}^{\prime}},$ $\quad\operatorname{rk}(\mathbf{s}\vee\mathbf{s}^{\prime})<\operatorname{rk}(\mathbf{s})+\operatorname{rk}(\mathbf{s}^{\prime})$
$c_{\mathbf{b}}c_{\mathbf{b}^{\prime}}\mathbf{y}_{\mathbf{s}}y_{\mathbf{s}^{\prime}}-c_{\mathbf{b}+\mathbf{b}^{\prime}}y_{\mathbf{s}\vee\mathbf{s}^{\prime}},$ $\mathbf{b},\mathbf{b}^{\prime}$ are bases of $\mathbf{s},\mathbf{s}^{\prime}$ so that $\mathbf{b}+\mathbf{b}^{\prime}$ is a basis for $\mathbf{s}\vee\mathbf{s}^{\prime}$.
We first show that $J\subset\ker\phi^{*}$. Suppose $\mathbf{s}$ and $\mathbf{s}^{\prime}$ are two combinatorial flats. If $\operatorname{rk}(\mathbf{s}\vee\mathbf{s}^{\prime})<\operatorname{rk}(\mathbf{s})+\operatorname{rk}(\mathbf{s}^{\prime})$, then the sum of any pair of bases of $\mathbf{s}$ and $\mathbf{s}^{\prime}$ is a dependent multiset, so $\phi(y_{\mathbf{s}}y_{\mathbf{s}^{\prime}})=0$. On the other hand, if $\operatorname{rk}(\mathbf{s}\vee\mathbf{s}^{\prime})=\operatorname{rk}(\mathbf{s})+\operatorname{rk}(\mathbf{s}^{\prime})$, then there are bases $\mathbf{b}$ and $\mathbf{b}^{\prime}$ such that $\mathbf{b}+\mathbf{b}^{\prime}$ is a basis for $\mathbf{s}\vee\mathbf{s}^{\prime}$, so $c_{\mathbf{b}}c_{\mathbf{b}^{\prime}}\phi(y_{\mathbf{s}}y_{\mathbf{s}^{\prime}})=c_{\mathbf{b}+\mathbf{b}^{\prime}}\phi(y_{\mathbf{s}\vee\mathbf{s}^{\prime}})$. Hence, $J\subset\ker\phi^{*}$.
Surjectivity of $\phi$ follows from the fact that $\phi(y_{\mathbf{s}})\smallfrown[Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}]$ is 1 if $\mathbf{s}=\mathbf{s}^{\prime}$ and 0 otherwise. Finally, it is evident from the relations that the degree $k$ part of $S/J$ has dimension at most $\dim H^{k}(Y_{V};\mathbb{Q})$, so $J=\ker\phi$. ∎
@@
In general, our proof of 1.1iii (Section 3.4) relies on high-powered results of [BHM+20b]. However, if $P$ is realized by $V\subset\mathbb{K}^{\mathbf{n}}$ with $\mathbb{K}$ of characteristic 0, then our work gives an independent proof of top-heaviness for $\mathcal{L}_{P}$ via arguments similar to those contained in [HW17, BE09].
@@
are injective when $2i\leq d$. These maps preserve the submodule $H^{*}(Y_{V})\subset I\!H^{*}(Y_{V})$, so they restrict to injections $H^{i}(Y_{V})\to H^{d-i}(Y_{V})$. 1.4 states that $\dim H^{k}(Y_{V})=|\mathcal{L}_{P}^{k}|$, so taking dimensions yields top-heaviness for $\mathcal{L}_{P}$.
@@ 5. Problems
We produce polymatroid Schubert varieties for polymatroids $P$ realizable over a field of characteristic 0. In this case, our work provides a geometric proof of the top-heavy property for $\mathcal{L}_{P}$. The proof of top-heaviness for realizable matroids [HW17] relies on being able to construct matroid Schubert varieties in positive characteristic.
@@
Construct polymatroid Schubert varieties over fields of positive characteristic. Use them to give a geometric proof of top-heaviness of $\mathcal{L}_{P}$ for any realizable polymatroid $P$.
@@
Matroid Schubert varieties are always normal, but this is untrue for polymatroid Schubert varieties.
@@
Determine when polymatroid Schubert varieties are normal.
@@
The ideals defining matroid Schubert varieties in $(\mathbb{P}^{1})^{N}$ are studied in [AB16], and found to be highly tractable: multidegrees, multigraded Betti numbers, and initial ideals can all be understood explicitly in terms of the combinatorics of the relevant matroid. We believe these objects will also admit combinatorial descriptions for polymatroid Schubert varieties. A major obstacle to verifying this is that we do not know the defining equations of polymatroid Schubert varieties.
@@
Find defining equations for polymatroid Schubert varieties.
Unlike those of matroid Schubert varieties, equations for polymatroid Schubert varieties cannot be obtained by mere homogenization of circuits, even in relatively simple cases.
@@
Let $V$ be the line in $\mathbb{C}^{3}$ spanned by $(1,1,1)$. A minimal set of polynomials defining $Y_{V}\subset\mathbb{P}^{3}$ is:
$b_{1}^{2}-2\,b_{0}b_{2}+4\,b_{0}b_{1}$
$b_{1}b_{2}-3\,b_{0}b_{3}+8\,b_{0}b_{2}-7\,b_{0}b_{1}$
$2\,b_{2}^{2}-3\,b_{1}b_{3}+12\,b_{0}b_{3}-46\,b_{0}b_{2}+56\,b_{0}b_{1}.$
The ideal generated by these polynomials is not the ideal of homogenized circuits of $V$. However, we observe that the number of minimal equations of a given multidegree $\mathbf{d}$ is equal to the number of circuits of $\widetilde{P}$ that represent $\mathbf{d}$. $\diamondsuit$
Along similar lines, [AB16, Theorem 1.5] gives a combinatorial formula for the multigraded Betti numbers of the ideal of a matroid Schubert variety.
@@
Find a combinatorial formula for the multigraded Betti numbers of the ideal of a polymatroid Schubert variety.
@@
The Chow class of a polymatroid Schubert variety $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ may not be multiplicity-free, and the coefficients $c_{\mathbf{b}}$ of $[Y_{V}]$ may depend on $\mathbf{n}$. We conjecture the following formula:
@@
If $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$, then the coefficient of $[\mathbb{P}^{\mathbf{b}}]$ in $[Y_{V}]$ is $c_{\mathbf{b}}=\binom{n_{1}}{b_{1}}\binom{n_{2}}{b_{2}}\cdots\binom{n_{N}}{b_{N}}$.
@@
The singularities of matroid Schubert varieties can be resolved in a “universal fashion” using the stellahdral variety of [BHM+20a].
@@
Resolve the singularities of polymatroid Schubert varieties. Is there a “universal” resolution?
@@
The $Z$-polynomial of a matroid $M$ is a combinatorial invariant defined in [PXY18]. When $M$ is realized by $V\subset\mathbb{K}^{N}$, the $Z$-polynomial is the Poincaré polynomial of the intersection cohomology of the matroid Schubert variety $Y_{V}$.
@@
Compute the intersection cohomology Betti numbers of a polymatroid Schubert variety. Are they combinatorially determined?

### v2

@@ 4. Schubert varieties of subspace arrangements
Proof of 1.7.
Let $S=\mathbb{Q}[y_{\mathbf{s}}:\text{$\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$}]$, graded by $\deg y_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. Using the notation of 4.16, define a map
@@
where $\mathbf{b}$ is any basis of $\mathbf{s}$. The image of $y_{\mathbf{s}}$ does not depend on $\mathbf{b}$ by 4.16.
@@
$y_{\mathbf{s}}y_{\mathbf{s}^{\prime}},$ $\quad\operatorname{rk}(\mathbf{s}\smile\mathbf{s}^{\prime})<\operatorname{rk}(\mathbf{s})+\operatorname{rk}(\mathbf{s}^{\prime})$
$c_{\mathbf{b}}c_{\mathbf{b}^{\prime}}y_{\mathbf{s}}y_{\mathbf{s}^{\prime}}-c_{\mathbf{b}+\mathbf{b}^{\prime}}y_{\mathbf{s}\smile\mathbf{s}^{\prime}},$ $\mathbf{b},\mathbf{b}^{\prime}$ are bases of $\mathbf{s},\mathbf{s}^{\prime}$ so that $\mathbf{b}+\mathbf{b}^{\prime}$ is a basis for $\mathbf{s}\smile\mathbf{s}^{\prime}$.
We first show that $J\subset\ker\phi$. Suppose $\mathbf{s}$ and $\mathbf{s}^{\prime}$ are two combinatorial flats. If $\operatorname{rk}(\mathbf{s}\smile\mathbf{s}^{\prime})<\operatorname{rk}(\mathbf{s})+\operatorname{rk}(\mathbf{s}^{\prime})$, then the sum of any pair of bases of $\mathbf{s}$ and $\mathbf{s}^{\prime}$ is a dependent multiset, so $\phi(y_{\mathbf{s}}y_{\mathbf{s}^{\prime}})=0$. On the other hand, if $\operatorname{rk}(\mathbf{s}\smile\mathbf{s}^{\prime})=\operatorname{rk}(\mathbf{s})+\operatorname{rk}(\mathbf{s}^{\prime})$, then there are bases $\mathbf{b}$ and $\mathbf{b}^{\prime}$ such that $\mathbf{b}+\mathbf{b}^{\prime}$ is a basis for $\mathbf{s}\smile\mathbf{s}^{\prime}$, so $c_{\mathbf{b}}c_{\mathbf{b}^{\prime}}\phi(y_{\mathbf{s}}y_{\mathbf{s}^{\prime}})=c_{\mathbf{b}+\mathbf{b}^{\prime}}\phi(y_{\mathbf{s}\smile\mathbf{s}^{\prime}})$. Hence, $J\subset\ker\phi$.
Surjectivity of $\phi$ follows from the fact that $\phi(y_{\mathbf{s}})\smallfrown[Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}]$ is 1 if $\mathbf{s}=\mathbf{s}^{\prime}$ and 0 otherwise. Finally, it is evident from the relations that the $y_{\mathbf{s}}$ generate $S/J$ as a $\mathbb{Q}$-vector space, so the degree $k$ part of $S/J$ has dimension at most $\dim H^{k}(Y_{V};\mathbb{Q})$, and therefore $J=\ker\phi$. ∎
@@
In general, our proof of 1.1iv (Section 3.4) relies on high-powered results of [BHM+20]. However, if $P$ is realized by $V\subset\mathbb{K}^{\mathbf{n}}$ with $\mathbb{K}$ of characteristic 0, then our work gives an independent proof of top-heaviness for $\mathcal{L}_{P}$ via arguments similar to those contained in [HW17, BE09].
@@
are injective when $i\leq d/2$. These maps preserve the submodule $H^{*}(Y_{V})\subset I\!H^{*}(Y_{V})$, so they restrict to injections $H^{i}(Y_{V})\to H^{d-i}(Y_{V})$. 1.7 states that $\dim H^{k}(Y_{V})=|\mathcal{L}_{P}^{k}|$, so taking dimensions yields top-heaviness for $\mathcal{L}_{P}$.
@@ 5. Problems
We produce Schubert varieties for polymatroids $P$ realizable over a field of characteristic 0. In this case, our work provides a geometric proof of the top-heavy property for $\mathcal{L}_{P}$. The proof of top-heaviness for realizable matroids [HW17] relies on being able to construct Schubert varieties of hyperplane arrangements in positive characteristic.
@@
Given a polymatroid $P$ realizable over a field of positive characteristic, construct a variety with an affine paving whose poset of strata if the lattice of combinatorial flats, thereby proving top-heaviness of $\mathcal{L}_{P}$.
@@
Schubert varieties of hyperplane arrangements are always normal, but this is untrue for Schubert varieties of subspace arrangements.
@@
Determine when Schubert varieties of subspace arrangements are normal.
@@
The ideals defining Schubert varieties of hyperplane arrangements in $(\mathbb{P}^{1})^{N}$ are studied in [AB16], and found to be highly tractable: multidegrees, multigraded Betti numbers, and initial ideals can all be understood explicitly in terms of the combinatorics of the relevant matroid. We believe these objects will also admit combinatorial descriptions for Schubert varieties of subspace arrangements. A major obstacle to verifying this is that we do not know their defining equations.
@@
Find defining equations for Schubert varieties of subspace arrangements.
Unlike those of Schubert varieties of hyperplane arrangements, equations for Schubert varieties of subspace arrangements cannot be obtained by mere homogenization of circuits, even in relatively simple cases.
@@
Let $V$ be the line in $\mathbb{C}^{3}$ spanned by $(1,1,1)$, which represents the polymatroid $P$ of rank 1 on one element. A minimal set of polynomials defining $Y_{V}\subset\mathbb{P}^{3}$ is:
$6\,x_{0}b_{12}+4\,x_{1}b_{12}+2\,b_{12}^{2}-6\,x_{0}x_{3}-3\,x_{1}x_{3}$
$x_{1}^{2}-2\,x_{1}b_{12}-2\,b_{12}^{2}+3\,x_{1}x_{3}$
$3\,x_{0}x_{1}+5\,x_{1}b_{12}+4\,b_{12}^{2}-3\,x_{0}x_{3}-6\,x_{1}x_{3}.$
The ideal generated by these polynomials is not the ideal of homogenized circuits of $V$. However, observe that there are three minimal equations of degree 2, and that the matroid lift $\widetilde{P}$ with respect to $\mathbf{n}=(3)$ is a uniform matroid of rank 1 on three elements, which has three circuits of size 2.
In general, we observe in examples that if $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$, then the number of minimal equations of a given multidegree $\mathbf{d}$ is equal to the number of circuits of the lift $\widetilde{P}$ with respect to $\mathbf{n}$ that represent $\mathbf{d}$. $\diamondsuit$
Along similar lines, [AB16, Theorem 1.5] gives a combinatorial formula for the multigraded Betti numbers of the ideal of a Schubert variety of hyperplane arrangement.
@@
Find a combinatorial formula for the multigraded Betti numbers of the ideal of a Schubert variety of subspace arrangement.
@@
The Chow class of a Schubert variety of subspace arrangement $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ may not be multiplicity-free, and the coefficients $c_{\mathbf{b}}$ of $[Y_{V}]$ may depend on $\mathbf{n}$. We conjecture the following formula:
@@
If $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$, then when $\mathbf{b}$ is a basis, the coefficient of $[\mathbb{P}^{\mathbf{b}}]$ in $[Y_{V}]$ is $c_{\mathbf{b}}=\binom{n_{1}}{b_{1}}\binom{n_{2}}{b_{2}}\cdots\binom{n_{N}}{b_{N}}$.
@@
The singularities of Schubert varieties of hyperplane arrangements can be resolved in a “universal fashion” using the stellahdral variety of [BHM+22].
@@
Resolve the singularities of Schubert varieties of subspace arrangements. Is there a “universal” resolution?
@@
The $Z$-polynomial of a matroid $M$ is a combinatorial invariant defined in [PXY18]. When $M$ is realized by $V\subset\mathbb{K}^{N}$, the $Z$-polynomial is the Poincaré polynomial of the intersection cohomology of the Schubert variety of hyperplane arrangement $Y_{V}$.
@@
Compute the intersection cohomology Betti numbers of a Schubert variety of subspace arrangement. Are they combinatorially determined?